OpenStudy (anonymous):

A shopkeeper allows rebate of 25% to buyer. He sells only smuggled goods ans as a bribe he pays 10% of the CP. If his CP=2500, then what should be the marked price if he desire to make profit of 9.09%?

OpenStudy (anonymous):

ok so it's like CP+bribe=total cost of the shopkeeper, right?

OpenStudy (anonymous):

=2500+10%(2500)=2500+250=2750

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

ok so now it says he wants 9.09% profit on that =2750+(9.09% of 2750)=3000

OpenStudy (anonymous):

ok i get it but answer is 4000? (

OpenStudy (anonymous):

wait it's not yet complete...so this 3000 he has got after rebate of 25% so to calculate that we should go in reverse order... that's increasing 3000 by 33.3%=4000(Ans) hope it's clear!

OpenStudy (anonymous):

Ah! ok thanks @aryandecoolest

OpenStudy (anonymous):

well..anytime )